Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-209_2001-215

Mission Operations Days: 2001/209 to 2001/215
July 28, 2001 00:00:00 GMT to August 03, 2001 00:00:00 GMT


MODIS Band 5 Detectors 1-20 (all) Gains Changed On 2001/212. See below for further details.


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On; #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 2 set of patches
  • FR Error Counting and Statistics (2 patches)
  • FR Band 5 Gain Change (20 commands)
Nominal
Inhibit Ids SetNoneNominal
TMON 66 (SRCA A)DisabledNominal
TMON 67 (SRCA B)DisabledNominal
TMON 68 (FIFO)EnabledNominal
TMON 69 (EPS Current)DisabledNominal

This Week's Completed MODIS Activities:

Saturday, July 28, 2001

None

Sunday, July 29, 2001

None

Monday, July 30, 2001

None

Tuesday, July 31, 2001

2001/212 13:49:40-13:57:16 RT Changing of Band 5 Gains to reflect 25% lower values

Wednesday, August 1, 2001

2001/213 21:33:32-21:45:00 ATC OA-15: SD/SDSM Open
2001/213 23:12:26-23:23:54 ATC OA-16: SD/SDSM Screened

Thursday, August 2, 2001

2001/214 00:22:28 RT Set SMIR Itwk/Vdet 79/110
2001/214 00:28:00-00:48:00 ATC 20 Extra Minutes FR Day Rate over Ground Track Night
2001/214 05:51:19-01:02:47 ATC OA-15: SD/SDSM Open
2001/214 01:11:59 RT Set SMIR Itwk/Vdet 79/190

Friday, August 3, 2001

None

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Friday, August 3, 2001

None

Saturday, August 4, 2001

None

Sunday, August 5, 2001

None

Monday, August 6, 2001

None

Tuesday, August 7, 2001

None

Wednesday, August 8, 2001

220/21:40:08-21:51:36 ATC OA-15: SD/SDSM Open
220/22:31:32-23:05:05 ATC OA-19: Full Radiometric
220/23:19:20-23:30:30 ATC OA-16: SD/SDSM Screened

Thursday, August 9, 2001

221/01:12:45-01:15:16 ATC OA-27: PV ECAL
221/02:50:39-02:55:10 ATC OA-28: PC ECAL

221/03:00:14 RT FOT Set IRU to high rate for roll maneuver; less pointing accuracy
221/05:12:23 ATC Move Formatter to Night ahead by 3:23 to ensure 50/50 day/night
221/05:30:14 ATC FOT begins spacecraft roll slew to -7.4496 degrees
221/05:34:27 ATC Set Formatter to Day Rate
221/05:34:29 ATC Set SCIABNORM Flag to ABNORM
221/05:34:31 ATC Set PC band DC Restore to OFF
221/05:34:33 ATC Set PV band DC Restore to OFF
221/05:34:35 ATC Set Sector Rotation to -3072 (Earth Aperture to view Space view and OBCs)
221/05:41:05 ATC Set Sector Rotation to Zero (Normal)
221/05:41:07 ATC Set PV band DC Restore to ON
221/05:41:09 ATC Set PC band DC Restore to ON
221/05:41:11 ATC Set SCIABNORM Flag to NORM
221/05:41:13 ATC Set Formatter to Night Rate
221/05:45:02 ATC Spacecraft has returned to nominal attitude
221/06:08:37 ATC Move Formatter to Day later by 3:23 to ensure 50/50 day/night
221/07:15:02 RT FOT set IRU to low rate after roll maneuver; nominal pointing

221/20:02:23 ATC Move Formatter to Night ahead by 3:23 to ensure 50/50 day/night
221/20:??:?? RT Set SMIR Itwk/Vdet 79/110
221/20:23:29 ATC Set Formatter to Day Rate
221/20:23:31 ATC Set SCIABNORM Flag to ABNORM
221/20:23:33 ATC Set PC band DC Restore to OFF
221/20:23:35 ATC Set PV band DC Restore to OFF
221/20:23:37 ATC Set Sector Rotation to -3072 (Earth Aperture to view Space view and OBCs)
221/20:30:07 ATC Set Sector Rotation to Zero (Normal)
221/20:30:09 ATC Set PV band DC Restore to ON
221/20:30:11 ATC Set PC band DC Restore to ON
221/20:30:13 ATC Set SCIABNORM Flag to NORM
221/20:30:15 ATC Set Formatter to Night Rate
221/20:??:?? RT Set SMIR Itwk/Vdet 79/190
221/20:58:37 ATC Move Formatter to Day later by 3:23 to ensure 50/50 day/night

Friday, August 10, 2001

None

Maneuvers:

The next MODIS Lunar Roll Calibration is scheduled for August 9, 2001 ( DOY 221 ) from 05:24:35 to 05:31:05. A second lunar calibration will be performed without a roll maneuver on the same day from 20:23:37 to 20:30:07. During the second roll, the SMIR Itwk/Vdet will be temporarily set to 79/110.

MODIS Anomalies:

The MODIS PS2 Shutdown Anomaly report was presented on Monday, July 30, 2001.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

MODIS Telemetry Trends:

Trending has resumed with the instrument on the A-side. All appears nominal at this time.

Non-MODIS Significant Events:

None.

Limited Life Item Status:

SRCA 10W Lamp #1: 189.0 of 500 hours
SRCA 10W Lamp #2: 144.7 of 500 hours
SRCA 10W Lamp #3: 153.9 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 557.9 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1667 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements